Astrid1:

2.The amount Lin’s sister earns at her part-time job is proportional to the number of hours she works. She earns$9.60 per hour. a.Write an equation in the form to describe this situation, where x represents the hours she works and represents the dollars she earns.

Dixieland:

do it ur self

JammarWalker:

Ill help

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dixieland do it ur self \(\color{#0cbb34}{\text{End of Quote}}\) I'm Just asking for help, Jeez.

Dixieland:

-.-

JammarWalker:

Do you know how to set up an equation?

Dixieland:

sorry

hhanan:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dixieland do it ur self \(\color{#0cbb34}{\text{End of Quote}}\) dont comment if your not going to help:)

Dixieland:

i mean i will if u need some

Astrid1:

I am not sure if I know how to set up the equation.

Dixieland:

okay

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Dixieland okay \(\color{#0cbb34}{\text{End of Quote}}\) leave im helping

Dixieland:

rude

Dixieland:

i was gonna help

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 I am not sure if I know how to set up the equation. \(\color{#0cbb34}{\text{End of Quote}}\) y=mx+b

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 I am not sure if I know how to set up the equation. \(\color{#0cbb34}{\text{End of Quote}}\) y=mx+b \(\color{#0cbb34}{\text{End of Quote}}\) y is a function of x, because for each input x there is only one output y.

JammarWalker:

The equation would be.........y=9.60x

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker The equation would be.........y=9.60x \(\color{#0cbb34}{\text{End of Quote}}\) where y is the amount Lin's sister earns and x is the number of hours she works

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker The equation would be.........y=9.60x \(\color{#0cbb34}{\text{End of Quote}}\) where y is the amount Lin's sister earns and x is the number of hours she works \(\color{#0cbb34}{\text{End of Quote}}\) HOPE THIS HELPS!!!!! 😁😁😁

Astrid1:

Okay,, so Would I have to multiply \[9.60\times x\]

Dixieland:

0.0

Dixieland:

just how

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 Okay,, so Would I have to multiply \[9.60\times x\] \(\color{#0cbb34}{\text{End of Quote}}\) NO all you needed was the equation right

Astrid1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @Astrid1 Okay,, so Would I have to multiply \[9.60\times x\] \(\color{#0cbb34}{\text{End of Quote}}\) NO all you needed was the equation right \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, i think...

Astrid1:

Well, Thanks!

Dixieland:

-.-

Dixieland:

i d k the awser good luck